OpenStudy (anonymous):

integrate cos(x)^0.5/(x)^0.5

OpenStudy (jamesj):

You're looking for this? \[\int\limits \sqrt{\frac{\cos x}{x}} \ dx\]

OpenStudy (jamesj):

Because if you are, there is no combination of standard, analytic functions which is equal to this integral.